OpenStudy (anonymous):

Find two vectors U and V. Given: U+V=<4,0> U is parallel to <0,-1> V is perpendicular to <0,-1> U=? V=?

OpenStudy (anonymous):

I have a question...

OpenStudy (anonymous):

Do you understand the dot product?

OpenStudy (anonymous):

A little, we just went over that yesterday in class but this is supposedly homework for the section before the dot product so I don't believe its suppose to be used here

OpenStudy (anonymous):

Okay, well, notice that U and V are perpendicular.

OpenStudy (anonymous):

Since U is parallel to \(\langle 0,-1\rangle \), it is in the same direction. Thus anything perpendicular to \(\langle 0,-1\rangle \) is perpendicular to U.

OpenStudy (anonymous):

This means that \[ u\cdot v = 0 \]

OpenStudy (anonymous):

This means that: \[ u_1v_1 + u_2v_2 = 0 \]

OpenStudy (anonymous):

We also know that \[ u_1+v_1 = 4 \]and \[ u_2+v_2 = 0 \]

OpenStudy (anonymous):

Oh wait, one more thing! \[ cu_1 = 0 \\ cu_2 = -1 \]If they are parallel, then the components are proportional!

OpenStudy (anonymous):

Okay, so how exactly do I put that all together to get U and V?

OpenStudy (anonymous):

Well \[ \mathbf u = \langle u_1,u_2\rangle \]

OpenStudy (anonymous):

\[ cu_1 = 0 \implies u_1 = 0 \]

OpenStudy (anonymous):

\[ u_1 +v_1 = 4\implies 0+v_1 = 4 \implies v_1=4 \]

OpenStudy (anonymous):

\[ \mathbf u\cdot \mathbf v = 0 \implies u_1v_1 + u_2v_2 = 0 \implies (0)(4) + u_2v_2=0\implies u_2v_2 = 0 \]

OpenStudy (anonymous):

\[ u_2+v_2 = 0\implies u_2 = -v_2 \]

OpenStudy (anonymous):

\[ u_2v_2=0\implies u_2(-u_2) = 0\implies -u_2^2=0\implies u_2 = 0 \]

OpenStudy (anonymous):

Wait. something isn't right here.

OpenStudy (anonymous):

I seem to be getting \(\mathbf u = \mathbf 0 = \langle 0,0\rangle\) and \(\mathbf v = \langle 4,0\rangle\). I suppose that makes sense. It's just a bit weird.

OpenStudy (anonymous):

I'm not sure you can say that \(\mathbf 0\) is parallel to any vector.

OpenStudy (anonymous):

I typed it into my webwork and it said its correct. must just be one of those weird problems. my professor makes up a lot of his own problems so it might have just been a weird one that he didn't think through

OpenStudy (anonymous):

The thing is, for all vectors \(\mathbf 0 \cdot \mathbf v = 0\).

OpenStudy (anonymous):

So technically \(\mathbf 0\) is perpendicular to all vectors.

OpenStudy (anonymous):

Also, \(c \cdot \mathbf 0 = \mathbf 0\), so \(\mathbf 0\) can only be said to be perpendicular to itself.

OpenStudy (anonymous):

I suppose if you let \(c =0\), then for all vectors \(c\mathbf v = \mathbf 0\).